OpenStudy (anonymous):

A certain quantity grows exponentially over time. The initial quantity at t = 0 is 1,000. The quantity grows by a factor of 3%. What is the quantity at t = 5 to the nearest whole number? need help please :) A. 116 B. 1126 C.1159 D. 3713

OpenStudy (anonymous):

1000*(1.03) = 1030 at t = 1 1000*(1.03) = 1060.9 at t = 2 1000*(1.03) = 1092.727 at t = 3 1000*(1.03) = 1125.508 at t = 4 1000*(1.03) = 1159.274 at t = 5

OpenStudy (anonymous):

ahhh okay gotcha. thanks